OpenStudy (anonymous):

solve the following system x-y=10 and x+y=8

OpenStudy (anonymous):

Multiple solutions.. any other info or directions?

OpenStudy (hoblos):

x-y=10 x+y=8 add the two equations 2x = 18 x=9 x+y=8 9+y=8 so y=-1

OpenStudy (anonymous):

Thanks!!!!

OpenStudy (hoblos):

u r welcome XD